A box contains six balls, each a different color. In how many orders can four balls be selected from the box?
Found 2 solutions by ikleyn, mathmate:
Answer by ikleyn(53765)

.
In 6*5*4*3 orders.


-------------------------
The words "In how many orders . . . " in the question assume and exactly mean that the ORDER IS IMPORTANT.

This notice relates to the other tutor solution.



Answer by mathmate(429)

Question:
A box contains six balls, each a different color. In how many orders can four balls be selected from the box?

Solution:
If order is not important, then number of ways
=C(6,4)=6!/(4!2!)=15
If order is important, then number of ways
=P(6,4)=6!/((6-2)!)=360